# Figure 1: #2

### v1

@@ 4.2 Result & Analysis
Table 3: The ablation study for introducing different reference masks.
4.3 Ablation Study
We conduct ablation study with various settings. Tab. 3 shows the situations for various reference guidance. Tab. 4 shows the improvement with the seperated dataset. Fig. 6 shows the influence about the quality of the reference mask and the size of the reference target.
Reference Guidance.
Providing the positive reference mask based on the reference image is the most intuitive way for users. Compared to the baseline, the positive guidance brings significant improvement. In PartImageNet, the improvement for NoC exceeds 0.5 clicks, which is large for IS. The IoU metric improvement of the first click more intuitively reflects the performance of the positive guidance. On the three datasets, the performance improvements for the SBD-based version it brings are 13.85%, 6.47%, 5.81%, respectively. All of these clearly indicate that the positive reference mask can greatly reduce ambiguity and improve the initial segmentation results. The negative reference mask is originally proposed to assist the positive one, but we also provide the performance of only the negative reference mask here. The improvement brought by the negative reference mask is not inferior to that of the positive reference mask at all. This is also easy to understand, because the negative masks can suppress the targets that users do not want to segment, and the generation of interactive ambiguity is often due to excessive segmentation of certain parts. We input both masks into RefCut and find that the performance is further improved. But the value of this improvement is very limited. This also indicates that some of the functions of positive and negative reference masks actually have a common impact. Therefore, if the number of images that need to be annotated is not too large, users can choose only the positive reference mask. In Tab. 4, we observe that for combined evaluation, RefCut achieve effective improvements in performance across all subsets. This demonstrates the applicability of RefCut to target diversity.
# NoC @80/85 Part (7162) Dual (4980) Whole (2353) Combine (14495)
$\clubsuit$ Baseline 3.62/4.80 3.28/4.18 1.44/1.69 3.15/4.08
RefCut 2.79/3.98 2.40/3.21 1.19/1.40 2.41/3.31
Table 4: Seperated improvement in PartImageNet [27]
Reference Mask Quality.
We gradually degrade the reference masks into polygons by interval sampling along the boundary points and connecting them. We adopt these masks as the reference and record the IoU&1 in Fig. 6 (a). We can see that when the mask quality is not too poor, RefCut can always achieve good performance. This is also in line with expectations, as we use the average feature of the reference mask as the prompt, so the impact of contour segmentation quality on the final performance is relatively small. This indicates that in practical use, users only need to provide a rough reference mask to achieve good results.
Reference Target Size.
We keep reducing the size of the reference image and the mask to get the performance of RefCut for the reference target at different scale ratios. We adopt these changed reference images and masks and record the IoU score with the first click in Fig. 6 (b). The results are as expected, the reduction in the size of the reference target leads to a decrease in the performance for RefCut. This also matches the results on the PASCAL-Part [12] dataset, which has a large number of very small targets, and using them as reference clearly does not yield good results.

### v2

@@ 4.3 Ablation Study
Table 3: The ablation study about NoC@X and IoU&X for introducing different reference masks.
Reference Mask Quality. We gradually degrade the reference masks into polygons by interval sampling along the boundary points and connecting them. We adopt these masks as the reference and record the IoU&1 in Fig. 5 (a). We can see that when the mask quality is not too poor, RefCut can always achieve good performance. This is also in line with expectations, as we use the average feature of the reference mask as the prompt, so the impact of contour segmentation quality on the final performance is relatively small. This indicates that in practical use, users only need to provide a rough reference mask to achieve good results.
Reference Target Size. We keep reducing the size of the reference image and the mask to get the performance of RefCut for the reference target at different scale ratios. We adopt these changed reference images and masks and record the IoU score with the first click in Fig. 5 (b). The results are as expected, the reduction in the size of the reference target leads to a decrease in the performance for RefCut. This also matches the results on the PASCAL-Part [43] dataset, which has a large number of very small targets, and using them as reference clearly does not yield good results.